An asset purchased on January 1, 2015 for $66,000 that has an estimated life of 10 years will have a book value on December 31, 2018 of:
Answer #1

Depreciation per year = 66,000/10 = 6,600 per year

Depreciation from Jan 1, 2015 to Dec 31,2018 = 6,600*4 years

= 26,400

Book value = Cost - Accumulated Depreciation

= 66,000 - 26,400

= 39,600